The perimeter of a rectangle is 310 feet. The length is 25 feet more than the width. Find the width and the length of the rectangle and the corresponding area of the rectangle
Answer by monika_p(71)

a - lenght
b- widht
Given
P=310
a= b+25
perimeter formula of rectangle
P=2a+2b
substitute for P and a
310=2*(b+25)+2b
b=65 then a= 90
Now you can find area using formula A=a*b